Commissioning: 22 May 2010 UT

Observer: John H, Doug M, Kate R (in Tucson)
Telescope Operator: SAllanson, WWack (in Tucson)
Telescope Support: Brian (LBTO)
Instrument Support: TShih (LBTO)
Adaptive Secondary Operator: Runa (LBTO)

Summary

The seeing was 0.7" to 0.8" for the first couple hours, but we were not able to collect any dat due to several hardware and software problems. By the time we were ready to collect data the see varied between 1.0" and 1.5", which continued through the rest of the night.

We remeasured the rotator center of the IRTC, in poor seeing, so the new value may not be very accurate.

We collected a set of field data in poor seeing (0.9"-1.5"), which may not be of any use.

Tested some of Doug's IDL routines to collimate.

Preparations

Details

2:45 Open Chamber. Problem with Left Shutter Door, required controllers to be restarted

2:50 pointto M5_0444 TelMode=TRACK PAM- Failed: timed

3:05 Left Shutter door open

3:17 pointto Arcturus TelMode=TRACK PAM- Success

3:20 Found star after ptspiral search. IE=-18.0 CA=-125. IA=-400

3:40 To fix coma changed CA= -27, IA= -400. Changed IE= -22

3:43 Finding BS9153 to collimate, lots of astigmatism.

3:56 Collimated

4:04 Changed guide star; have to change parameters on the AGW (John Hill doing this).

4:06 Sent preset again with off-axis guide star # 10, collimated.

4:14 Going to BS9152 to see brighter central star

4:16 IRTC is not cooling, cannot see any stars in view.

4:27 Tim, Runa, and Steve going up to look at the cooling

4:32 The IRTC is unable to mount /newdata, attempting to fix here before writing IssueTrack

4:47 They power cycled the IRTC camera to make cooling reboot. Had no affect; now rebooting the Windows control computer. Downtime ~30minutes now.

04:50 Powered back up, IRTC is now cooling.

05:07 IRTC guide star found, on-axis star is out of focus.

05:15 Attempting to fix /newdata access problem with Dave Thompson contacted via phone.

05:22 Archive machine did not have Samba running after reboot a few nights ago. Dave turned on Samba server on Archive1. Connection to /newdata restored.

05:21 Cannot write to /newdata, Dave called again.

05:38 Fixed rights to /newdata, can now write to archive. Downtime ~17 minutes

05:44 setFocus right (previous at 14.2) changed to 15; to 18; to 20; to 25.

05:49 Determined 25mm will be the best value; thus, a delta of 10mm.

06:04 Going to M5 star 1416

06:12 Putting rotator to zero to do a rotator center point test @ RFBG.

06:20 takeseq @ .05 secs 80 frames @ 2Hz, only got 10 images. Accidentally put in wrong frame rate and exposure time. IRTC files set: 061444

06:23 takeseq @ .02 secs, 100 frames @ 10 Hz, not what we wanted. IRTC files set: 062237

06:26 Returning to active mode, to collimate again.

06:42 Rotation cube #8 is the useful cube with a 300 degree arc.

06:44 Slewing to BS9177, guide star zero, to get better seeing.

06:46 Seeing is bad at this elevation too ~ 2"; waiting for seeing to improve to potentially take field measurements.

06:56 Slewing to BS9166, guide star 15: due to wind direction and better elevation

06:59 Guide star has lots of coma, collimated it out.

07:27 Seeing has been really bad, just started to get ~1". Going to slew to a lower elevation star to try to take field data. Down time ~ 27 mins.

07:28 Slewing to BS9177 guide star 41.

07:37 Constructing a script to take field data on BS9177

07:42 Seeing in improving ~.9" to 1"

08:06 Starting the script on BS9177 called field_BS9177.cl to collect field data, seeing is ~.9" to 1.1"

08:07 Killed script; Needed to reset the pointing model, star was on edge of acquisition image, need to realign star to center.

08:11 Changed IE= -25 and CA= -12.

08:22 Collimation is bad, seeing getting worse as well, temperature dropping, humidity rising.

8:23 Preset sent BS9177, guide star 41.

8:37

GStar wfscimage#
Sequence 1
0 3 - 6
40 7 - 9
0 10 - 12
57 13 - 15
0 16 - 18
46 19 - 21
0 22 - 24
13 25 - 27
0 28 - 30
30 31 - 33
0 34 - 36
54 37 - 39
0 40 - 42
32 43 - 45

09:20 Guiding stopped unexpectedly. ERROR: failed to request full guide cam readout (after sending a preset to BS9177 guide star 41)

09:24 Stopping GCS right, and restarting, restarting worked. Apparently GCS got stuck when talking to camera. GCS Issue Track submitted #2748.

GStar wfscimage#
Sequence 2
0 46 - 49
14 50 - 52
0 53 - 55
41 56 - 58
0 59 - 61
57 62 - 64
0 65 - 67
46 68 - 70
0 71 - 73
13 74 - 76
0 77 - 79
30* 80* - 82*

10:11 *Noticed the guide star image for 82 was up to the left from center, 83 had the same issue.

10:16 *Found script error- missed a RFBG in field_BS9177A.cl- images #80 - #88 are all rubbish.

10:23 Restarted IRAF script on image # 89. Seeing was oscillating up to 2.5", has since improved to 1.1"- 1.6".

GStar wfscimage#
Sequence 3
0 89 - 92
30 93 - 95
0 96 - 98
50 99 - 101
0 102 - 104

10:42 Finished RFBG Field data

10:50 pointto BS9177 guide star 0

11:10 Ran two IDL scripts iif_preset and iif_offset, but worked properly after modification

11:13 Running IDL script collimate_once, many errors encountered.

11:30 Fixed errors, rerunning collimate_once, working well this time.

11:34 Ran collimate_multiple, did not read out to telescope, fixed the error.

11:42 Collimate_multiple script lost guiding, no star in field after secondary mirror focus value was sent.

11:44 Sent iif_preset and guide star missing (but preset) in acquisition image. Getting too bright.

11:50 Shutting down for the night.

-- DougMiller - 22 May 2010
Topic revision: r6 - 24 May 2010, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback